OpenStudy (anonymous):

Find three solutions of the equation. y=-5x+1 a. (–2, 11), (1, –4), (0, 1) c. (–2, 11), (1, –4), (2, –8) b. (0, 1), (3, –16), (–2, 11) d. (1, –4), (0, 1), (–1, 5)

OpenStudy (anonymous):

what about this one On which of the following lines does the point (–8, 31) lie?

OpenStudy (anonymous):

m=-5, x=-8, y=31 y-y1=m(x-x1) y-31=-5(x-(-8)) y-31=-5x-40 y=-5x-9